Two of the lines of the atomic hydrogen spectrum have wavelengths of 656?nm and 410?nm. If these fall at normal incidence on a grating with 8200 slits/cm,what will be the angular separation of the two wavelengths in the first-order spectrum?
Answer #1

apply d sin theta = mL

here d = 1/8200 = 1.219 um

for wave length 1, L1 = 656 nm

sin theta = 1* 656nm/1.219 um = 0.538

theta = 32.55 deg

for L2, sin theta = 1* 410nm/1.219 um

sin theta = 0.336

theta = 19.633 deg

angular separation is   32.55-19.633 = 12.91 deg